OpenStudy (anonymous):

A gumball machine contains 6 lemon,4 lime, 3 cherry and 2 orange gumballs. 5 gumballs are obtained. a. How many different sets of 5 gumballs are possible? b. How many will contain 2 lemon and 3 lime gumballs? c. Find the probability that the 5 gumballs dispensed by the machine include: 1. 2 lemon and 3 lime 2. 3 cherry and 2 orange 3. 2 lemon, 2 lime, and 1 orange 4. Only lemon 5. Only lime 6. No lemon

OpenStudy (anonymous):

I don't think any one will answer this unless they have a life

OpenStudy (anonymous):

Always worth a try- and you were trying to say is "I don't think anyone will answer this unless they don't have a life".

OpenStudy (anonymous):

A. 29 sets of 5 gumballs

OpenStudy (anonymous):

errr can you show the work or explain for a,b, and c1+2?

OpenStudy (anonymous):

Work for A. First you figure out the possible outcomes so 6*4=24*3=72*2=144 if there are 5 possible outcomes than you do 144/5 = 28.8 so you round up to 29 possible and that is the correct answer (I also looked it up online after I figured it out for myself.

OpenStudy (anonymous):

B. 1/29 Work for B. there is only one way to get 2 lemon and 3 lime only one set of ways out of 29 so it is 1/29 or just 1

OpenStudy (anonymous):

And i have lots of hw to do and tests to study for so the other people on this problem need to help you!